size : 10,12,14,16,18,20
frequency: 3,7,?,20,8,5
mean is 15 find the missing frequency

Answer by math_tutor2020(3833)


x = the missing frequency value

Given data
Size101214161820
Frequency37x2085


Or we can format the table like this
SizeFrequency
103
127
14x
1620
188
205


Multiply each size with its paired frequency. Then add up the results.
10*3 + 12*7 + 14*x + 16*20 + 18*8 + 20*5 = 14x+678

This is then divided over the total frequency (3+7+x+20+8+5 = x+43) to determine the mean.

mean = (14x+678)/(x+43) = 15
(14x+678)/(x+43) = 15
14x+678 = 15(x+43)
14x+678 = 15x+645
678-645 = 15x-14x
33 = x
x = 33


Answer: 33